### 03 Jan 2024 v1

# DDN-SLAM: Real-time Dense Dynamic Neural Implicit SLAM with Joint Semantic Encoding

Abstract
We propose DDN-SLAM, a real-time dense neural implicit semantic SLAM system designed for dynamic scenes. While existing neural implicit SLAM systems perform well in static scenes, they often encounter challenges in real-world environments with dynamic interferences, leading to ineffective tracking and mapping. DDN-SLAM utilizes the priors provided by the deep semantic system, combined with conditional probability fields, for segmentation.By constructing depth-guided static masks and employing joint multi-resolution hashing encoding, we ensure fast hole filling and high-quality mapping while mitigating the effects of dynamic information interference. To enhance tracking robustness, we utilize sparse feature points validated with optical flow and keyframes, enabling loop closure detection and global bundle optimization. Furthermore, DDN-SLAM supports monocular, stereo, and RGB-D inputs, operating robustly at a frequency of 20-30Hz. Extensive experiments on 6 virtual/real datasets demonstrate that our method outperforms state-of-the-art approaches in both dynamic and static scenes.
1 Introduction
In the fields of robotics and virtual reality, the Neural Radiance Fields (NeRF) [70, 67, 66, 65, 71, 72] have shown tremendous potential. There is a significant focus on real-time dense SLAM (Simultaneous Localization and Mapping) systems [50, 39, 60, 19, 21, 47] to combine with NeRF. Dense reconstruction [65, 68, 73] is crucial for scene perception and understanding, providing improvements over sparse reconstruction [13, 14, 58, 6, 9]. While traditional SLAM systems have achieved remarkable results in tracking and sparse reconstruction [10, 20], low-resolution and discontinuous surface reconstruction fails to meet the requirements [7, 11, 46]. Recently, neural implicit SLAM methods [42, 44, 18, 17, 16] have demonstrated excellent performance, with systems capable of real-time reconstruction [23, 24, 43] and inference in large-scale virtual/real scenes. These SLAM systems outperform traditional SLAM methods in terms of texture details, memory consumption, noise handling, and outlier processing.
Although current neural implicit SLAM systems have achieved good reconstruction results in static scenes [24, 78, 8, 40], many real-world environments are often affected by dynamic objects, especially in applications such as robotics or autonomous driving, which involve complex physical environments and may also have low-texture areas or significant changes in lighting and viewing angles. Current neural implicit SLAM systems are unable to achieve effective tracking and reliable reconstruction in such environments. The depth information and pixel interference caused by dynamic objects lead to inaccurate tracking results, while severe occlusions result in distorted reconstructions and ghosting artifacts. In addition, compared to traditional SLAM systems [1, 4, 2, 3, 5], neural implicit SLAM often lacks accurate loop closure detection and global bundle adjustment, which leads to inevitable drift accumulation and ultimately affects the quality of reconstruction.
To tackle these challenges, we propose DDN-SLAM. It can utilize prior semantic information to segment static and dynamic feature points in the scene. Additionally, it constructs sparse flow constraints for potential dynamic points and constrains keyframes. In terms of reconstruction, we propose a skip voxel strategy based background filling strategy that is more cost-effective than traditional methods. We utilize the semantic mask joint multi-resolution hash encoding [25] to achieve a balance between real-time performance, low memory consumption, and high-quality geometry and texture details. It also eliminates interference from dynamic objects and restricts the generation of ghosting artifacts through reasonable pixel filling. For tracking, we utilize feature points validated by semantic and flow threads to construct a complete bundle adjustment and loop closure detection framework, which achieves more robust tracking results in challenging scenes. In summary, our contributions can be summarized as follows:
  • We propose DDN-SLAM, a semantic SLAM system that enables dense reconstruction in real dynamic scenes. Its robust semantic system effectively removes dynamic interference, enabling real-time tracking and high-quality mapping in dynamic scenes.
  • We propose a framework which combines flow consistency checks, enabling robust tracking in challenging scenes. It supports real-time loop closure detection, bundle adjustment in dynamic or static scenes, and camera inputs from monocular, stereo, and RGB-D configurations.
  • We conduct extensive experiments on six static/dynamic datasets and perform various tests in challenging scenes. Our approach achieves effective tracking and reconstruction in scenes where current neural implicit SLAM methods struggle. Moreover, it delivers competitive tracking and reconstruction results in static scenes.
2 Related Work
2.1 Traditional Dynamic Visual SLAM
In most SLAM systems [33, 39, 50, 51], eliminating false data from dynamic objects [52, 53] that are commonly present in the real world is an important problem. It holds significant implications for both camera tracking and dense reconstruction [59, 60, 62]. Early researchers often used reliable constraints or multi-view geometric methods for feature matching, but they often failed to accurately segment subtly moving objects. Additionally, there was a lack of viable solutions for handling potential moving objects. Methods relying on optical flow consistency [54, 55, 56, 57] could differentiate by detecting inconsistent optical flow generated by moving objects, but with lower accuracy.
In recent years, deep object detection networks that incorporate semantic information have been employed for pixel-level segmentation to more accurately deal with dynamic objects in real-world environments [22]. However, the depth models used in traditional SLAM methods often incur high computational costs, impacting real-time performance. Moreover, traditional methods suffer from an inability to perform scene representation and reasonable hole filling for complex textures and geometric details due to discrete or incomplete surface structures.
2.2 NERF-BASED SLAM
The achievements of NeRF [25, 26, 83, 28, 29, 30, 66] in 3D scene representation [32, 34, 35, 36, 37, 38] have attracted widespread attention. In recent work, neural implicit SLAM systems and similar work [61, 63] have shown impressive performance in high-quality rendering and accurate 3D reconstruction. iMAP [23] proposes a real-time updating system using a single MLP decoder, and NICE-SLAM [42] extends the system to achieve larger scene representation, addressing the scene limitations of iMAP [23]. ESLAM [18] uses multi-scale vertically aligned feature planes for representation and decodes compact features directly into TSDF. Co-SLAM [44] adopts a joint encoding method based on coordinates and one-blob to balance geometric accuracy and inference speed. GO-SLAM [31] introduces an effective loop closure strategy that enables global correction but may impact tracking speed and efficiency. However, these methods result in severe errors in reconstruction and tracking in dynamic scenes due to interference from high/low dynamic objects. That leads to incorrect data association and cumulative drift amplification in pose estimation. The dynamic objects as interference information affecting the reconstruction results, causing artifacts or catastrophic forgetting issues. Although NICE-SLAM [42] proposes a pixel filtering method based on depth threshold to achieve effective tracking and reconstruction in low dynamic scenes, its inference speed and tracking accuracy decrease significantly in high dynamic scenes with occlusion and interference. Meanwhile, in challenging scenarios [77], neural implicit SLAM systems often get weak in pose estimation, resulting in tracking loss or significant pose drift. Neural implicit SLAM methods usually cannot match traditional methods in tracking accuracy due to the lack of loop detection and global BA. To tackle these limitations, we use deep learning-based depth segmentation and optical flow anomaly detection to obtain static encoding. Our system supports keyframe-based loop detection and global bundle adjustment, achieving competitive performance in both dynamic and static datasets compared to baseline methods.
3 Method
3.1 Joint Semantic Coordinate and Hash Encoding
In order to address the interference introduced by dynamic objects in the real world, a feature point distribution model is adopted. We combine statistics and background segmentation to obtain a static mask and eliminate the influence of dynamic feature points [80]. We utilize YOLOv5 [81] to obtain dynamic bounding box priors and form a set containing dynamic points. This allows us to avoid computing the fundamental matrix in traditional VSLAM methods [82] and instead identify dynamic points using a combination of depth and anomalous optical flow.
Since points outside the bounding boxes can be considered as static points following a Gaussian distribution, we use the median of all depth values as the baseline dynamic point threshold for each bounding box. Depth values below this threshold are considered as foreground, with the foreground depth following a Gaussian distribution. We use the T-distribution test to detect outliers, and a point’s T-value $\delta T$ is defined (only for keyframes that satisfy the threshold)
$\delta T=\left(\frac{d-\frac{1}{n}\sum_{k=1}^{n}d_{k}}{\sqrt{\frac{\sum_{j=1}^{n}\left(d_{j}-\frac{1}{n}\sum_{k=1}^{n}d_{k}\right)^{2}}{n-1}}}\right)^{2}$ (1)
where $d$ represents the depth of a point. We set the threshold as $T_{th}$. Points with chi-square values greater than the threshold are considered background points. To further filter dynamic feature points, we introduce sparse optical flow validation. The matched points obtained from sparse optical flow are from frame $m$ and $n$, represented as $P_{m}$ and $P_{n}$. We build a Gaussian distribution model using static points and use a t-distribution test to filter dynamic points. The covariance matrix of the coordinates of the points is defined as
$\Sigma=\left[\begin{array}[]{cc}\left(s^{2}\right)^{n}&0\\ 0&\left(s^{2}\right)^{n}\end{array}\right].$ (2)
We set $s$ to 2 and $n$ to 3. $P_{m}$ is chosen as the starting point, and $P_{n}$ is chosen as the ending point. The vector between two matching points is defined as $V$. We set $N$ to represent the number of selected static points. We apply filtering on the background points outside the bounding box to obtain $V_{mn}^{\text{static}}$. The optical flow vectors here also follow a Gaussian distribution $G(x;\mu,\sigma)$, where $\mu$ is the mean and $\sigma$ is the standard deviation of the data. We remove the optical flow of feature points whose magnitudes are greater than $3\sigma$. This is done to prevent the influence of abnormal optical flow points caused by photometric changes on the computation. We set the average optical flow of each point as $V_{s}^{\text{static}}=\frac{1}{n}\sum_{m,n}^{N}V_{mn}^{\text{static}}$. The t-distribution value $\delta V$ between two matched points is defined as:
$\delta V=(V-V_{s}^{\text{static}})^{T}\Sigma^{-1}(V-V_{s}^{\text{static}})$ (3)
where degrees of freedom $F$ is $n-1$. The threshold is set to $V_{th}$. If more than 10% of points within a box are dynamic, we consider the bounding box to contain a dynamic object. This ensures the consistency of object motion in the scene. Foreground points within dynamic boxes are considered as dynamic points and excluded during pose initialization.
Figure 1: Our method achieves accurate segmentation of multiple targets while removing unreliable feature points and preserving static feature points within the bounding box in high dynamic scenes of the Bonn [85] dataset. The red markers indicate the feature points recovered by using the method described in Sec. 3.3.
Figure 2: System Overview. Our DDN-SLAM system consists of two main modules: tracking and mapping, divided into four threads. Our four threads optimize alternately. The segmentation thread detects and segments dynamic feature points and pixels, suppressing potential feature points. The tracking thread extracts feature points, receives feature points filtered by conditional filtering for tracking, obtains static optical flow, generates keyframes and camera poses. The mapping thread receives background segmentation masks for high and low dynamics, performs keyframe generation and volume rendering. The loop detection thread detects loops and performs global bundle adjustment. Our system can be updated in real time.
3.2 Volume Rendering and Semantic Encoding
For representing 3D scenes in the presence of occlusion, we employ a volumetric rendering process combined with multi-resolution hash encoding [25]. This allows for effective geometric detail representation and adaptability to large-scale scenes[83], where each corner of a unit cell corresponds to a hashed grid that embeds encoding features. The given position $x$ can be mapped to the respective positions on each resolution grid, and trilinear interpolation is performed on the grid cell corners to obtain concatenated encoding feature vectors $\boldsymbol{\gamma}_{t}(\mathbf{x}_{t},\mathbf{t})\in\mathbb{R}^{c}$ After feature concatenation, we input them into an MLP. More specifically, our SDF network $F_{\Phi_{sdf}}$ is composed of a single multi-layer perceptron (MLP), with learnable parameters $\Phi_{sdf}$. It takes the point position $x$ and the corresponding hash encoding $h_{\Phi_{hash}}(\mathbf{x})$ as inputs and predicts the signed distance field (SDF) as:
$\Phi_{sdf},\mathbf{t}=F_{\Phi_{sdf}}(\mathbf{x},h_{\Phi_{hash}}(\mathbf{x})).$ (4)
To estimate the color $\mathbf{\Phi}(\mathbf{x})$, we utilize a color network $F_{\Phi_{\text{color}}}$ that processes the learned geometric feature vector $\mathbf{t}$, position $\mathbf{x}$, and SDF gradient $\mathbf{n}$ with respect to $\mathbf{x}$:
$\mathbf{\Phi}(\mathbf{x})={F_{\Phi_{\text{color}}}(\boldsymbol{\gamma}(\mathbf{x}),{t})}.$ (5)
We adopt hierarchical sampling to obtain $N_{\text{strat}}$ samples on each ray, and additionally sample $N_{\text{imp}}$ points near the surface. For all $N=N_{\text{strat}}+N_{\text{imp}}$ points on the ray. We sampled N and denoted as $x_{k}=\mathbf{o}+t_{k}\mathbf{d}$. This process represents the projection of a ray $\mathbf{r}$ from the camera center $\mathbf{o}$ through a pixel along its normalized viewing direction $\mathbf{d}$. We adopt the SDF representation and express its opacity as $\alpha_{i}$.
$\alpha_{k}=\max\left(\frac{\Omega\left(\Phi_{s}\left(\mathbf{x}_{k}\right)\right)-\Omega\left(\Phi_{s}\left(\mathbf{x}_{k+1}\right)\right)}{\Omega\left(\Phi_{s}\left(\mathbf{x}_{k}\right)\right)},0\right)$ (6)
where $\Omega$ is the sigmoid function. $T_{k}=\prod_{j=1}^{k-1}\left(1-\alpha_{j}\right)$ denotes the cumulative transmission rate, which represents the proportion of light reaching the camera. Similar to [25], we utilize skip voxel strategy. We maintain a series of K multi-scale occupancy grids. The occupancy status of each grid is continuously traversed using a Digital Differential Analyzer (DDA)[12] after 16 training iterations, and grid updates are performed. To improve efficiency, we employ a ray propagation sampling method to reduce unnecessary sampling. When $T_{k}$ drops below a threshold $\varepsilon$, we stop ray propagation and set the remaining contribution to 0. With the weight ${T_{k}}$ We use MLP predictions to render the color and depth for each ray:
$\hat{\mathbf{C}}=\sum_{k=1}^{N}{T_{k}}{\alpha_{k}}\mathbf{\Phi}(\mathbf{x})\quad\hat{\mathbf{D}}=\sum_{k=1}^{N}{T_{k}}{\alpha_{k}}{t_{k}}$ (7)
For mapping, we adopt an input keyframe-based approach. Inspired by [1][76][84], we adaptively update the background mask to enhance mapping quality. Unlike traditional dynamic SLAM methods which require complete traversal and completion of all input frames. We only need to perform pixel control on selected keyframes and specific dynamic pixels. We retained static pixels that do not interfere with the tracking process to ensure that our method can correctly distinguish between dynamic and static information without interference.
We obtain the sparse optical flow vectors $V_{s}^{\text{static}}$ for the filtered static feature points mentioned in Sec. 3.1, along with the pixel mask $M_{d}$ in the bounding box. Due to the varying motion speeds of different parts of dynamic objects, additional artifacts are introduced in the form of residual interference. Because the jump sampling strategy filter out low occupancy rate pixels, we can further refine $M_{d}$ by removing low dynamic pixels to reduce its occupancy rate. Based on temporal consistency, we calculate the optical flow vectors $M_{si}^{\text{dynamic}}$ between two consecutive frames for each pixel in $M_{d}$. (In this paper, we use keyframes involved in mapping and their corresponding adjacent frames as reference frames for calculation.) The L1 norm of the difference is obtained as:
$G_{t}=\frac{1}{N}\sum_{i=1}^{N}\left|V_{s}^{\text{static}}-M_{si}^{\text{dynamic}}\right|.$ (8)
We obtain a difference sequence $\{G_{s1},G_{s2},\ldots,G_{sn}\}$. Assuming the sequence follows a normal distribution, we set the threshold as $G_{t}$. When $G_{si}<G_{t}\ \text{{and}}\ 1<G_{si}$, we set the remaining contribution of the corresponding pixel’s depth to 0. To ensure real-time performance and effective removal of low dynamic pixels. If the removed pixels exceed 60% of the total dynamic pixels or the keyframe sampling rate ($n/N$) is less than 60%. We generate a new keyframe from the keyframe list to prevent a decrease in mapping quality due to excessive removal. For the correctness of backend optimization and loop detection, the additionally generated keyframes are not involved in bundle optimization and loop detection.
To supervise the network, we use the color loss $\mathcal{L}_{rgb}$ between the input image $c$ and the rendered image $\hat{c}$ and depth loss. The color and depth rendering loss are set as:
$\mathcal{L}_{rgb}=\frac{1}{N}\sum_{n=1}^{N}\left(\hat{\mathbf{C}}_{s}-\mathbf{C}_{s}\right)^{2},\mathcal{L}_{d}=\frac{1}{\left|R_{s}\right|}\sum_{r\in R_{s}}\left(\hat{\mathbf{D}}_{r}-{\mathbf{D}}_{s}\right)^{2}$ (9)
where $\hat{\mathbf{D}}_{r}$ denotes the filtered ray set with valid depth, and ${\mathbf{D}}_{s}$ corresponds to the filtered pixel depth set. $R_{s}$ is the batch of rays with depth contribution. Specifically, in the monocular mode, we utilize the predicted depth obtained from [75] as supervision. Integrating uncertainty-aware loss, we define the Mapping loss as:
$\mathcal{L}_{M}=\lambda_{r}\mathcal{L}_{\text{rgb }}+\lambda_{d}\mathcal{L}_{\mathrm{D}}.$ (10)
3.3 Tracking and Bundle Adjustment
Camera Tracking. In the presence of ground truth depth, we perform tracking using a system based on [74]. In monocular mode, we utilize the prior generated by [75] to obtain the scale factor $M$. Using the pose matrix $P_{e}$ obtained from the essential matrix, we can express $M$ as:
$d_{c}^{\prime}=K\cdot T\cdot d_{w},\quad M=\frac{1}{N}\sum_{i=1}^{N}\frac{d_{c}^{i}}{d_{p}^{i}}$ (11)
where $d_{c}^{\prime}$ is the feature point in camera coordinate system, $K$ is the camera intrinsic matrix, $T$ is the pose matrix, and $d_{w}$ is the position of the feature point in the world coordinate system. $d_{c}^{i}$ is the predicted depth of the feature point, and $d_{p}^{i}$ is the predicted depth from the network. The corrected pose matrix $P_{c}$ is obtained as $P_{c}=M*P_{e}$.
Loop Detection. In order to achieve effective tracking in large-scale scenes, we introduce loop detection to correct cumulative errors. We retrieve the three most similar keyframes from the keyframe groups using the DBoW2 with a similarity threshold of 0.8. The matching points here refer to the filtered feature points mentioned in Sec3.3. We check for geometric consistency. If at least three out of the 5 co-observed keyframes (already in the map) satisfy the condition of successful matching with the candidate keyframe group, we consider a common area is detected. Once the location recognition is successful, we align the active keyframe $K_{a}$, co-observed keyframe $K_{m}$, and all map points through transformation.
Figure 3: Reconstruction results on the dynamic sequence of TUM RGB-D [49]. The reconstruction results of the RGB-D image stream in high dynamic scenes are shown in the black boxes. The reconstruction results of the RGB-D image stream in low dynamic scenes are shown in the blue boxes. The reconstruction results of the monocular mode in high dynamic sequence are shown in the red boxes. Our method is capable of effectively reconstructing and removing occlusion interference, thus restoring the scale of monocular images.
Completion Rate walking/xyz walking/half walking/static walking/rpy sitting-xyz sitting-half AVG
LC-CRF SLAM [54] 100% 0.020 0.030 0.028 0.051 0.013 0.026 0.028
Crowd-SLAM [55] 100% 0.019 0.036 0.007 0.044 0.017 0.025 0.027
ORB-SLAM2 [10] 93% 0.721 0.462 0.386 0.785 0.009 0.026 0.373
NICE-SLAM [42] 79% 0.420 0.732 0.491 N/A 0.029 0.134 0.362
ESLAM [60] 61% 0.432 N/A 0.075 N/A 0.034 0.063 0.151
Co-SLAM [44] 44% 0.714 N/A 0.499 N/A 0.065 0.023 0.325
DDN-SLAM(RGB) 100% 0.028 0.041 0.025 0.089 0.013 0.031 0.037
DDN-SLAM 100% 0.014 0.023 0.010 0.039 0.010 0.017 0.020
Table 1: The tracking results on 8 dynamic sequences in TUM RGB-D [49](ATE[m]). We also present the average tracking Completion Rate of 6 scenes. The best results are indicated in bold, and the second-best results are indicated with a horizontal line. N/A indicates failure in all 5 measurements and the total number of frames with successful tracking is less than 50% of the total. Our method achieves more stable tracking results compared to traditional methods, while maintaining a running speed of 20fps.
Bundle Adjustment. During pose optimization, we should avoid including dynamic feature points in the optimization process. Therefore, through the filtering process mentioned in Sec. 3.1, we assign weights $w_{i}$ to static points and dynamic points as follows:
$w_{i}=1-{G_{a}}\left(\delta T,T_{\text{th }}\right)$ (12)
where $\delta T$ is the t-statistic of the point and $T_{th}$ is the threshold mentioned in Sec. 3.1. $G_{a}$ is the Gaussian function used to describe the closeness between the point and the background more accurately. We update the weights of each map point based on the corresponding feature points in the current frame. We set the definition of the new joint optimization cost function for the reprojection error as $e_{i}(x)$. To avoid excluding some potential static points, we introduce mean reprojection error $e_{m}(x)$ as a check before BA. It is defined as:
$e_{i}(x)=w_{i}\left|x_{i}-\pi\left(RX_{i}+T_{e}\right)\right|$ (13)
$e_{m}(x)=\frac{1}{N}\sum_{k=1}^{N}\left|x-\pi\left(R_{k}X+T_{ek}\right)\right|$ (14)
where $\pi(\cdot)$ is the projection transformation from the camera coordinate system to the plane coordinate system, and $\rho$ is the robust Huber loss function. $T_{e}$ as the translation vector. We compute the t-value ($T_{em}$) for the average reprojection error to recover deleted feature points, similar to the application of conditional probability for $e_{m}$ as in 3.1. We set the threshold as $T_{eth}$. In particular, we disable the function of eliminating potential feature points in static scenes. In the pose estimation, we jointly optimize the rotation matrix $R$, translation vector $T_{e}$, Mapping of 3D map points to image coordinates $P_{e}$ and the weights $w_{i}$ of each map point. We obtain the minimum reprojection error $L_{rp}$ as follows:
$\min_{\left\{[R\mid T_{e}]\right\},\left\{P_{e}\right\}}L_{rp}=\underset{\left\{[R\mid T_{e}]\right\},\left\{P_{e}\right\}}{\operatorname{argmin}}\sum_{i=0}^{N}\rho\left(e_{i}(x)^{T}\mathbf{\Theta}e_{i}(x)\right)$ (15)
where $\mathbf{\Theta}$ is the information matrix related to the scale of the feature points. We performed BA to optimize the the keyframes and camera poses and the observed map points in these keyframes. The optimized keyframes and mapped points were then passed to NeRF for further processing.
4 Experiment
4.1 Experimental Setup
Implementation Details. Our system is implemented on a system with a NVIDIA RTX 3090ti GPU and an Intel Core i7-12700K CPU operating at 3.60 GHz. We use the YOLOv5 network pre-trained on the COCO dataset[69] as the backbone of our object detection system. Due to the page limit the parameter settings and detailed ablation experiments can be found in the supplementary materials.
Dataset. We evaluated our system on a total of 2 dynamic datasets, 1 challenging dataset with large viewpoints and weak textures, and 3 static datasets. The datasets used for evaluation include TUM RGB-D [49], Bonn [85], OpenLORIS-Scene [77], Replica[48], ScanNet[45], and EuRoC [41]. The TUM RGB-D [49] dataset consists of 6 high-dynamic sequences, 2 low-dynamic sequences, and 3 static sequences. The Bonn [85] dataset includes 6 high-dynamic range sequences and 2 sequences with complex occlusions. The challenging scenes dataset comprises 7 indoor sequences from the OpenLORIS-Scene [77] dataset, demonstrating large viewpoint changes and lighting variations. The static datasets consist of 6 sequences from the ScanNet [45] dataset, 8 sequences from the high-quality virtual scene dataset Replica[48], and 3 Sequences on the EuRoC [41] Dataset. Due to the page limit, we include more experimental results in the supplemental material.
Completion Rate balloon balloon2 move move2 crowd crowd2 person person2 AVG
LC-CRF SLAM [54] 100% 0.027 0.024 0.079 0.186 0.027 0.098 0.046 0.041 0.066
Crowd-SLAM [55] 100% 0.037 0.014 0.026 0.029 0.019 0.035 0.046 0.099 0.038
ORB-SLAM2 [10] 91% 0.065 0.230 0.320 0.039 0.496 0.989 0.692 0.079 0.363
NICE-SLAM [42] 87% 2.442 2.018 0.177 0.832 1.934 3.582 0.245 0.536 1.470
ESLAM [60] 64% 0.203 0.235 0.190 0.129 0.416 1.142 0.845 7.441 1.325
Co-SLAM [44] 100% 0.211 0.480 0.076 0.200 51.983 2.019 0.695 0.758 7.052
DDN-SLAM 100% 0.018 0.041 0.020 0.032 0.018 0.023 0.043 0.038 0.029
Table 2: Tracking results on 8 dynamic sequences in the Bonn[85] dataset(ATE[m]). We also present the average tracking Completion Rate of 8 scenes. The best results are indicated in bold, and the second-best results are indicated with a horizontal line. For methods that fail to track, we report the best result obtained before the failure. Our method maintains stable operation at 15FPS while preserving the multi-object tracking capability, aided by TensorRT[79] acceleration.
Completion Rate office1-1 office1-2 office1-3 office1-4 office1-5 office1-6 office1-7 AVG
ORB-SLAM2 [10] 100%↑ 0.074 0.072 0.015 0.077 0.161 0.080 0.100 0.082
ElasticFusion [4] 100%↑ 0.112 0.118 0.022 0.100 0.144 0.113 0.192 0.114
DS-SLAM [20] 77%↑ 0.089 0.101 0.016 0.021 0.052 0.129 0.077 0.069
NICE-SLAM [42] 100%↑ 0.110 0.142 0.040 0.232 0.130 0.104 0.251 0.144
ESLAM [60] 100%↑ 0.101 0.152 0.010 0.235 0.151 0.101 0.201 0.150
Co-SLAM[44] 99%↑ 0.112 0.173 0.043 0.141 0.174 0.134 0.186 0.138
GO-SLAM [31] 100%↑ 0.120 0.160 0.033 0.183 0.942 0.092 0.142 0.239
DDN-SLAM 100%↓ 0.058 0.071 0.011 0.056 0.120 0.050 0.105 0.067
Table 3: Tracking results on 7 challenging sequences in the Openloris-Scene [77] dataset(RGB-D). We also present the average tracking Completion Rate of 7 scenes. The best results are highlighted in red, the second-best results are highlighted in green, and the third-best results are highlighted in yellow. The evaluation was performed using the computation tools provided by the authors.
Figure 4: Reconstruction results on the dynamic sequence of Bonn [85]. Our method is able to reconstruct stably in occluded environments, and compared to traditional methods [54], our approach has advantages in restoring texture details and inferring unknown regions.
Metrics and Baseline. We evaluate the reconstruction quality using Depth L1 (cm), Accuracy (cm), Completion (cm), Completion Ratio (%), and PSNR (dB). For assessing camera tracking quality in dynamic scenes, we use (ATE)RMSE. Our main baselines for reconstruction quality and camera tracking include iMAP[23], NICE-SLAM[42], ESLAM[60], GO-SLAM [31] and Co-SLAM[44]. We utilized the replication code provided by [78] to achieve better performance reproduction for iMAP[23]. For a fair comparison, we introduce classical traditional dynamic SLAM methods for tracking results in dynamic scenes, including LC-CRF SLAM [54], Crowd-SLAM [55], ElasticFusion [4], ORB-SLAM2 [10], and DS-SLAM [20].
4.2 Evaluation in Dynamic Scenes
TUM RGB-D [49] Dynamic Sequence Evaluation. To evaluate our system on TUM RGB-D [49] dynamic sequences, we compare it with LC-CRF SLAM [54], Crowd-SLAM [55], ORB-SLAM2 [10], NICE-SLAM [42], ESLAM [18] and Co-SLAM [44]. Our system achieves effective rendering, effectively avoids occlusion interference caused by high-dynamic human subjects, reasonably fills scene holes, and achieves high-precision geometric detail reconstruction with minimal artifacts, as shown in Figure 4. We have achieved tracking results that are competitive with traditional methods. It is important to highlight that existing neural implicit methods often face challenges in achieving effective tracking in high-dynamic scenes, as shown in Table 2. Due to severe accumulated drift, current neural implicit SLAM methods are unable to complete reliable mapping.
Bonn [85] Dataset Dynamic Sequence Evaluation. Our tracking method demonstrates robustness and achieves competitive tracking results, as demonstrated in Table 3. Compared to the TUM RGB-D [49] dataset, the Bonn [85] dataset exhibits more complex dynamic scenes, severe occlusions, and the presence of multiple moving objects simultaneously. The reconstruction results are shown in Fig. 5. In comparison to traditional dynamic SLAM methods, our approach achieves better results. The current neural implicit systems have shown significant drift or tracking failures in most scenarios.
4.3 Evaluation in Static Scenes
[…]
Table 4: Runtime analysis on the ScanNet [45] dataset. The best result is shown in bold, and the second-best result is indicated with an underline. The results of GO-SLAM [31] are from its paper, while the results of other methods are reported based on the average of 6 measurements. Our method demonstrates competitiveness in memory consumption and runtime speed, enabling it to run on lighter-weight platforms, which is a prerequisite for widespread adoption of neural implicit SLAM.
ScanNet [45] Dataset Experimental Results. We assessed the average running speed, memory usage, mapping time, and GPU memory usage (on a 3090ti GPU) across six scenes. The results indicate that our method achieves higher running speed and reduces memory consumption. (Please refer to the supplementary material for detailed tracking results.)
TUM RGB-D [49] Dataset Experimental Results. (including monocular and RGB-D ATE data and reconstruction result images) We tested the performance of monocular and RGB-D modes on this dataset. In three static sequences, our RGB-D mode achieved better tracking results than traditional methods, as shown in Table 4. It changed the disadvantage of neural implicit methods over traditional methods in this dataset. At the same time, our monocular mode also achieved stable tracking results, correctly recovering scale and pose.
Openloris-Scene [77] Dataset Experimental Results. We conducted reconstruction and tracking evaluation on the challenging Openloris-Scene [77] dataset. Our method is significantly better than existing neural implicit methods, and there is no significant drift as in NICE-SLAM [42] and Co-SLAM [44], as shown in Table 5. Thanks to the mature tracking system, our method is able to achieve accurate relocalization in extremely challenging scenes and improve the completion rate of tracking. Even under large viewpoint changes and low texture-low resolution image constraints, we still achieve high-quality reconstruction, which indicates that our 3D reconstruction results benefit from the joint encoding method and adaptive keyframe generation strategy.
[…]
Table 5: Tracking Results on TUM RGB-D [49] dataset. ATE RMSE [cm] (↓) is used as the evaluation metric. We compared neural implicit methods and traditional approaches, and our results achieved superior tracking performance over the traditional methods on the TUM RGB-D [49] dataset. Additionally, our monocular results ensured robustness and accurate scale recovery.
5 Conclusion
We propose DDN-SLAM, a neural implicit semantic SLAM system that enables stable tracking in complex occlusion and challenging environments. It achieves dense indoor scene reconstruction through 2D-3D scene understanding. Our system is more aligned with real-world requirements and can flexibly support various inputs while demonstrating superior performance. It accurately captures geometric features and texture details, while enabling long-term updates and global optimization. Our experiments demonstrate that our method effectively avoids interference from severe occlusion on mapping and tracking, and achieves state-of-the-art performance on various datasets. Limitations. The application of our system on mobile platforms such as robot platforms has promising potential. Furthermore, more reasonable depth estimation and detail optimization with spatio-temporal consistency may lead to better mapping results.